# EvalMuse-40K: A Reliable and Fine-Grained Benchmark with Comprehensive Human Annotations for Text-to-Image Generation Model Evaluation

### v1

@@ Abstract
Recently, Text-to-Image (T2I) generation models have achieved significant advancements. Correspondingly, many automated metrics have emerged to evaluate the image-text alignment capabilities of generative models. However, the performance comparison among these automated metrics is limited by existing small datasets. Additionally, these datasets lack the capacity to assess the performance of automated metrics at a fine-grained level. In this study, we contribute an EvalMuse-40K benchmark, gathering 40K image-text pairs with fine-grained human annotations for image-text alignment-related tasks. In the construction process, we employ various strategies such as balanced prompt sampling and data re-annotation to ensure the diversity and reliability of our benchmark. This allows us to comprehensively evaluate the effectiveness of image-text alignment metrics for T2I models. Meanwhile, we introduce two new methods to evaluate the image-text alignment capabilities of T2I models: FGA-BLIP2 which involves end-to-end fine-tuning of a vision-language model to produce fine-grained image-text alignment scores and PN-VQA which adopts a novel positive-negative VQA manner in VQA models for zero-shot fine-grained evaluation. Both methods achieve impressive performance in image-text alignment evaluations. We also use our methods to rank current AIGC models, in which the results can serve as a reference source for future study and promote the development of T2I generation. The data and code will be made publicly available.
@@ 1 Introduction
Recently, advanced Text-to-Image (T2I) models [24, 30, 10, 31, 33, 34, 47] are capable of generating numerous impressive images. However, these models may still generate images that fail to accurately match the input text, such as inconsistency in quantities [19, 44, 43]. Given the high cost and inefficiency of manual evaluation, developing a reliable automatic evaluation metric and corresponding benchmark is vital. They can effectively assess existing model performance and guide improvements for future models.
@@
Since traditional evaluation metrics, such as FID [13] and CLIPScores [12], are not well-suited for assessing the consistency of T2I model generation, recent works [16, 42, 45] have explored the construction of new evaluation metrics. These methods introduce diverse evaluation approaches leveraging various multi-modal models. For example, VQAScore [21] is constructed by asking the Visual Question Answering (VQA) model “Does this figure show {text}?” and obtaining the likelihood of answer “Yes” as the alignment score. However, such a simple VQA manner cannot handle the fine-grained matching problem well. TIFA [16] and VQ2 [45] set up related questions for multiple elements split from the prompt, then averaged the answers to produce a final alignment score. These methods allow for fine-grained evaluation of image-text alignment but are also limited by the performance of the VQA model. Moreover, the relationship between element-based alignment scores and overall human preferences has yet to be explored. Therefore, to better explore the performance of existing T2I evaluation methods, we contribute a new benchmark, EvalMuse-40K, featuring fine-grained human annotations of image-text pairs.
@@
EvalMuse-40K includes 4K prompts, 40K image-text pairs, and more than 1M fine-grained human annotations. To ensure the diversity of prompts, EvalMuse-40K includes 2K real prompts and 2K synthetic prompts, where the real prompts are sampled from DiffusionDB [41]. We categorize the real prompts from multiple dimensions and use the MILP [39] strategy to ensure the category balance of the final sampled prompts. The synthesized prompts are then constructed for specific skills in image-text alignment, such as quantity and location. By synthesizing specific prompts and sampling from real prompts, EvalMuse-40K is able to accurately evaluate specific model problems while also better exploring model performance in real scenarios. For fine-grained evaluation, we use a large language model for the elemental splitting of prompts and question generation, and to increase the diversity of the generated images, we generate images using a variety of T2I models. Compared with the previous benchmark (see Tab. 1 for details), EvalMuse-40K not only scores image-text alignment but also performs more fine-grained annotations for elements split from the prompts. Additionally, we annotate different types of structural problems that may appear in generated images. This comprehensive and fine-grained human evaluation in EvalMuse-40K enables an analysis of current automated T2I model evaluation metrics and their correlation with human preferences, ultimately supporting the improvement and development of the T2I evaluation system.
@@
In addition to the comprehensive benchmark, we introduce two methods, FGA-BLIP2 and PN-VQA, for image-text alignment evaluation. FGA-BLIP2 uses vision-language models to jointly fine-tune the image-text alignment scores and element-level annotations so that the models can output the overall scores and determine whether the generated images match the elements in the prompt. Additionally, we employ a variance-weighted optimization strategy to account for prompts’ ability to generate diverse images. PN-VQA constructs specific positive-negative question-answer pairs, allowing MLLMs to evaluate the fine-grained alignment of image-text pairs. With these two methods, we can obtain a scoring model aligned with human preferences and perform fine-grained alignment analysis and evaluation using MLLMs.
@@
Finally, we select some representative prompts to regenerate images and analyze existing AIGC models using our proposed evaluation methods. We will release our data, models, and code to support further study.
@@
  • EvalMuse-40K collects a large number of human annotations and ensures a more robust evaluation of automated metrics’ performance by using a more balanced prompt with a larger number of generative models.
  • EvalMuse-40K categorizes elements during fine-grained annotation, allowing us to evaluate the accuracy of current automated metrics in assessing specific skills at a fine-grained level.
@@
  • We introduce a new automated evaluation metric, PN-VQA, which employs positive-negative questioning to enhance correlation with human ratings.
@@ 2 Related Work
Figure 1: Overview of EvalMuse-40K, consisting of (a) data collection, (b) data annotation, and (c) evaluation methods. (a) We collected 2K real prompts and 2K synthetic prompts, using the MILP [39] sampling strategy to ensure category balance and diversity. Prompts are further divided into elements, and corresponding questions are generated. We also use various T2I models to generate images. (b) During data annotation, we labeled the fine-grained alignment levels of image-text pairs, structural problems about the generated images, and some extra information. For annotated data with large differences in overall alignment scores, we performed re-annotation to ensure the reliability of the benchmark. (c) We proposed two effective image-text alignment evaluation methods: one is FGA-BLIP2, using a fine-tuned vision-language model for direct fine-grained scoring of image-text pairs, and another is PN-VQA, adopting positive-negative VQA manner for zero-shot evaluation.
@@
Image-Text Alignment Benchmarks. Many different benchmarks have been proposed to evaluate the image-text alignment of T2I models. Early benchmarks are small-scale and mostly rely on captions from existing datasets like COCO  [7, 16, 27, 32], focusing on some sample skills. Other benchmarks (e.g., HPDv2 [43] and Pick-a-pic [19]) use side-by-side model comparisons to evaluate the quality of the generated images. Recently, benchmarks like DrawBench [34], PartiPrompt [47] and T2I-CompBench [17] have introduced a set of prompts and focused on evaluating specific skills of generative models, including counting, spatial relationships, and attribute binding. Moreover, some benchmarks (e.g., GenAI-Bench [20] and RichHF-18K [25]) provide human annotations on image-text align scores to validate the relevance of automated metrics to human preference. For fine-grained evaluation, benchmarks like TIFA [16] and SeeTRUE [45] extract elements from prompts and generate corresponding questions.
However, prompts in current benchmarks often focus on specific T2I model skills, which differ from prompts used by real users. The proposed EvalMuse-40K addresses this by containing 2K real prompts sampled from DiffusionDB [41] and 2K synthetic prompts for specific skills, enabling a more comprehensive evaluation of image-text alignment in image generation. Additionally, it provides large-scale human annotations on both overall and fine-grained aspects of image-text alignment, allowing deeper analysis of how well current evaluation metrics align with human preference.
Automated Metrics for Image-Text Alingment. Image generation models initially use Fréchet Inception Distance (FID) [13] to calculate the distributional differences between generated images and validation images. Besides, IS [35] and LPIPS [48] have been used to evaluate the quality of the generated images. However, since these metrics do not effectively evaluate image-text alignment, recent work primarily reports CLIPScore [12], which measures the cosine similarity of text and image’s features. With the strong performance of BLIP2 [23], BLIP2Score has also been adopted as an alignment metric similar to CLIPScore.
@@
Human preference models (e.g., ImageReward [44], PickScore [19], and VNLI [45]) improve assessment capabilities by using a vision-language model like CLIP, fine-tuned on large-scale human ratings. However, these models primarily rely on side-by-side image comparison annotations to learn human preferences, making it difficult to obtain accurate alignment scores and perform fine-grained evaluations. We address this by using a vision-language model to directly predict both overall and fine-grained alignment scores for image-text pairs.
For fine-grained evaluation, TIFA [16], VQ2 [45], and Gecko [42] split prompts into elements and generate questions to assess fine-grained issues in generated images using VQA models. However, these generated questions can sometimes diverge from the original prompt’s content. To address this, we design a new positive-negative question template that provides VQA models with more accurate prompt context, improving alignment with human preference.
@@ 3 EvalMuse-40K Benchmark
In this section, we detail EvalMuse-40K, a reliable and fine-grained benchmark with comprehensive human annotations for T2I evaluation. We present the overall architecture of EvalMuse-40K in Fig. 1.
In the construction process, we employ various strategies to ensure the diversity and reliability of our benchmark. Next, we will introduce the construction process of EvalMuse-40K from three aspects: prompt collection, image generation, and data annotation followed by statistical analysis of the data.
@@ 3.1 Prompts Collection
Real Prompts from Community. Most prompts used by current benchmarks are generated from templates with LLMs or manually crafted, with only a small portion originating from real-world users. This results in a gap between model evaluation and actual user needs. To address this, we sampled 2K real prompts from DiffusionDB [41], which contains 1.8M prompts specified by real users.
To be specific, we first randomly sampled 100K prompts from DiffusionDB as the source dataset. To ensure diversity in the final sampled prompts, we classified the prompts in four aspects: subject category, logical relationship, image style, and BERT [18] embedding score. The specific categories of each aspect are shown in the supplementary material. Except for the BERT embedding score, which is computed, the other aspects are categorized using GPT-4 [1], and prompts can belong to more than one category. After labeling the 100K prompts, we improved the data shaping method proposed in [39] by using Mixed Integer Linear Programming (MILP) to ensure that the sampled data is approximately uniformly distributed across each category. The specific sampling strategy and the distribution of the sampled data are shown in the supplementary material.
Synthetic Prompts for Various Skills. To comprehensively evaluate the generative models’ skills, we employed specific templates and a rich corpus to generate 2K synthetic prompts using GPT-4. These prompts were divided into six categories: (1) Object Count; (2) Object Color and Material; (3) Environment and Time Setting; (4) Object Activity and Perspective Attributes; (5) Text Rendering; and (6) Spatial Composition Attributes. For each category, we used different templates and GPT-4 to generate reasonable and natural prompts. The specific generation methods and details are provided in the supplementary material.
Element Splitting and Question Generation. To achieve fine-grained annotation and evaluation, we performed element splitting and question generation on the 4K collected prompts. In contrast to the word-level annotation used in RichHF [25] and Gecko [42], we split the prompt into fine-grained elements, which increases the consistency of the annotations across different humans. In addition, categorizing each element allows us to examine the model’s capabilities in various aspects at a fine-grained level. We followed the element categorization strategy from TIFA [16] but divided the question generation process into two parts: first splitting the prompts into elements, and then generating specific questions for each element. We generate only simple judgment questions with yes or no answers, making the question generation process more controllable and the evaluation more explainable. The generated questions are also filtered and regenerated to ensure that each element has a corresponding question. The templates used for element splitting and question generation are shown in the supplementary material.
@@ 3.2 Image Generation
To ensure the diversity of the generated images, we selected 20 different types and versions of diffusion-based generative models, considering the strong performance of diffusion models [14]. Meanwhile, to ensure the differentiation among the selected models, we classify the selected models into four groups, (1) basic stable diffusion [33] models, such as SD v1.2, SD v1.5, SD v2.1; (2) advanced open-source generative models, such as SDXL [31], SSD1B [11], SD3 [10], HunyuanDiT [24], Kolors [38], PixArt-$\alpha$ [6], PixArt-$\Sigma$ [4], IF [8], Kandinsky3 [2], Playground v2.5 [22]; (3) efficient generative models, such as SDXL-Turbo [36], LCM-SDXL, LCM-SSD1B [29], LCM-PixArt [5], SDXL-Lightning [26]; and (4) proprietary generative models, such as Dreamina [9] and Midjourney v6.1 [15]. For each prompt, we randomly sample a subset of models for image generation, applying default parameters to ensure the quality of the generated images. This process results in 40K image-text pairs being generated using 4K prompts, ensuring a diverse dataset for annotation.
@@ 3.3 Data Annotation
In this section, we describe how we performed the data annotation. We first define the content and templates of the annotations and then detail the entire annotation process.
Annotation Content and Templates. The annotation includes image-text alignment, structural problems, and extra information (see Fig. 1(b)). In terms of image-text alignment, the annotator first scores the alignment using a 5-point Likert scale, like TIFA [16]. Then, for the fine-grained elements in the prompt, the annotator needs to label whether they are aligned with the image or not. Annotators also mark structural issues in the generated images, categorized into three main types: humans, animals, and objects. For human figures, structural issues are further subdivided by specific regions, such as the face, hands, and limbs. To address potential inaccuracies in element splitting by GPT-4, we introduced a new splitting confidence label, enabling annotators to flag instances with incorrect splitting during annotation. Additionally, because some prompts originate from real users and may contain unclear meanings, we added a label to indicate whether a prompt is meaningful.
Annotation Process. To improve the quality of annotation, our annotation process is divided into three stages. 1) Pre-annotation: We formulate straightforward annotation standards to train the annotators, using a small amount of data for pre-annotation. The pre-annotated data are then carefully reviewed. Based on the issues encountered during the annotation and review process, we refine the evaluation standards to ensure greater clarity and consistency. 2) Formal annotation: During formal annotation, each image-text pair is labeled by three annotators, with an additional annotator assigned for quality control. Furthermore, the annotators will identify any NSFW content in the generated images and determine whether they should be discarded. 3) Re-annotation: For instances where alignment scores from the three annotators show significant discrepancies (range $\geq 2$), we conduct re-annotation to reduce subjective bias. Ultimately, each image-text pair is annotated by 3 to 6 annotators, and the average score is used as the final label.

### v2

@@ Abstract
Text-to-Image (T2I) generation models have achieved significant advancements. Correspondingly, many automated metrics emerge to evaluate the image-text alignment capabilities of generative models. However, the performance comparison among these automated metrics is constrained by the limited scale of existing datasets. Additionally, these datasets lack the capacity to assess the performance of automated metrics at a fine-grained level. In this study, we contribute an EvalMuse-40K benchmark, gathering 40K image-text pairs with fine-grained human annotations for image-text alignment-related tasks. In the construction process, we employ various strategies such as balanced prompt sampling and data re-annotation to ensure the diversity and reliability of our benchmark. This allows us to comprehensively evaluate the effectiveness of image-text alignment metrics for T2I models. Meanwhile, we introduce two new methods to evaluate the image-text alignment capabilities of T2I models: FGA-BLIP2 which involves end-to-end fine-tuning of a vision-language model to produce fine-grained image-text alignment scores and PN-VQA which adopts a novel positive-negative VQA manner in VQA models for zero-shot fine-grained evaluation. Both methods achieve impressive performance in image-text alignment evaluations. We also apply our methods to rank existing AIGC models, providing results that can serve as a reference for future research and foster the development of T2I generation.
@@ 1 Introduction
Recently, advanced Text-to-Image (T2I) models [24, 30, 10, 31, 33, 34, 47] are capable of generating numerous impressive images. However, these models may still generate images that fail to accurately match the input text, such as inconsistency in quantities [19, 44, 43]. Given the high cost and inefficiency of manual evaluation, developing a reliable automatic evaluation metric and corresponding benchmark is vital. They can effectively evaluate the performance of existing models and provide guidance for improvements in future models.
@@
Since traditional evaluation metrics, such as FID [13] and CLIPScores [12], are not well-suited for assessing the consistency of T2I models, recent works [16, 42, 45] have explored the construction of new evaluation metrics. These methods introduce diverse evaluation approaches leveraging various multi-modal models. For example, VQAScore [21] is constructed by asking the Visual Question Answering (VQA) model “Does this figure show {text}?” and obtaining the likelihood of answer “Yes” as the alignment score. However, such a simple VQA manner cannot handle the fine-grained matching problem well. TIFA [16] and VQ2 [45] decompose the prompt into multiple elements, formulate related questions for each, and then average the answers to generate a final alignment score. These methods allow for fine-grained evaluation of image-text alignment but are also limited by the performance of the VQA model. Moreover, the relationship between element-based alignment scores and overall human preferences has yet to be explored. Therefore, to better explore the performance of existing T2I evaluation methods, we contribute a new benchmark, EvalMuse-40K, featuring fine-grained human annotations of image-text pairs.
@@
EvalMuse-40K includes 4K prompts, 40K image-text pairs, and more than 1M fine-grained human annotations. To ensure the diversity of prompts, EvalMuse-40K includes 2K real prompts and 2K synthetic prompts, where the real prompts are sampled from DiffusionDB [41]. We categorize the real prompts from multiple dimensions and use the MILP [39] strategy to ensure the category balance of the final sampled prompts. The synthesized prompts are then constructed for specific skills in image-text alignment, such as quantity and location. By synthesizing specific prompts and sampling from real prompts, EvalMuse-40K is able to accurately evaluate specific model problems while also providing a more comprehensive assessment of model performance in real-world scenarios. For fine-grained evaluation, we use a large language model for the elemental splitting of prompts and question generation, and to increase the diversity of the generated images, we generate images using a variety of T2I models. Compared with the previous benchmark (see Tab. 1 for details), EvalMuse-40K not only scores image-text alignment but also performs more fine-grained annotations for elements split from the prompts. Additionally, we annotate different types of structural problems that may appear in generated images. This comprehensive and fine-grained human evaluation in EvalMuse-40K enables an analysis of current automated T2I model evaluation metrics and their correlation with human preferences, ultimately supporting the improvement and development of the T2I evaluation system.
@@
In addition to the comprehensive benchmark, we introduce two methods, FGA-BLIP2 and PN-VQA, for image-text alignment evaluation. FGA-BLIP2 uses vision-language models to jointly fine-tune the image-text alignment scores and element-level annotations so that the models can output the overall scores and determine whether the generated images match the elements in the prompt. Additionally, we employ a variance-weighted optimization strategy to account for prompts’ ability to generate diverse images. PN-VQA constructs specific positive-negative question-answer pairs, allowing MLLMs to evaluate the fine-grained alignment of image-text pairs. With these two methods, we can obtain a scoring model aligned with human preferences and perform fine-grained alignment analysis and evaluation.
@@
Finally, we select some representative prompts to regenerate images and analyze existing AIGC models using our proposed evaluation methods. Our data, models, code, and results serve as valuable resources to support further research.
@@
  • EvalMuse-40K collects a large number of human annotations and ensures a more robust evaluation of automated metrics by using a well-balanced set of prompts and a diverse range of generative models.
  • EvalMuse-40K categorizes elements during fine-grained annotation, enabling the evaluation of existing automated metrics’ accuracy in assessing specific skills at a fine-grained level.
@@
  • We introduce a new automated metric, PN-VQA, which employs positive-negative questioning to enhance the correlation with human ratings.
@@ 2 Related Work
Figure 1: Overview construction of EvalMuse-40K, consisting of (a) data collection, (b) data annotation, and (c) evaluation methods. (a) We collect 2K real-world prompts and 2K synthetic prompts, using the MILP [39] sampling strategy to ensure category balance and diversity. Prompts are further divided into elements, and corresponding questions are generated. We also use various T2I models to generate images. (b) During data annotation, we label the fine-grained alignment levels of image-text pairs, structural problems of the generated images, and some extra information. For annotated data with large differences in overall alignment scores, we perform re-annotation to ensure the reliability of the benchmark. (c) We propose two effective image-text alignment evaluation methods: one is FGA-BLIP2, using a fine-tuned vision-language model for direct fine-grained scoring of image-text pairs, and another is PN-VQA, adopting positive-negative VQA manner for zero-shot evaluation.
@@
Image-Text Alignment Benchmarks. Many different benchmarks have been proposed to evaluate the image-text alignment of T2I models. Early benchmarks are small-scale and mostly rely on captions from existing datasets like COCO  [7, 16, 27, 32], focusing on limited sample skills. Other benchmarks (e.g., HPDv2 [43] and Pick-a-pic [19]) use side-by-side model comparisons to evaluate the quality of the generated images. Recently, benchmarks like DrawBench [34], PartiPrompt [47], and T2I-CompBench [17] have introduced a set of prompts and focused on evaluating specific skills of generative models, including counting, spatial relationships, and attribute binding. Moreover, some benchmarks (e.g., GenAI-Bench [20] and RichHF-18K [25]) provide human annotations on image-text align scores to validate the relevance of automated metrics with human preference. For fine-grained evaluation, benchmarks like TIFA [16] and SeeTRUE [45] extract elements from prompts and generate corresponding questions. However, prompts in current benchmarks often focus on specific T2I model skills, which differ from prompts used by real users. The proposed EvalMuse-40K addresses this by containing 2K real prompts sampled from DiffusionDB [41] and 2K synthetic prompts for specific skills, enabling a more comprehensive evaluation of image-text alignment in image generation. Additionally, it provides large-scale human annotations on both overall and fine-grained aspects of image-text alignment, allowing deeper analysis of how well current evaluation metrics align with human preference.
@@
Automated Metrics for Image-Text Alignment. Image generation models initially use FID [13] to calculate the distributional differences between generated images and validation images. Besides, IS [35] and LPIPS [48] have been used to evaluate the quality of generated images. However, since these metrics do not effectively evaluate image-text alignment, recent work primarily reports CLIPScore [12], which measures the cosine similarity of text and image’s features. With the strong performance of BLIP2 [23], BLIP2Score has also been adopted as an alignment metric similar to CLIPScore. Human preference models (e.g., ImageReward [44], PickScore [19], and VNLI [45]) improve assessment capabilities by using a vision-language model like CLIP, fine-tuned on large-scale human ratings. However, these models primarily rely on side-by-side image comparison annotations to learn human preferences, making it difficult to obtain accurate alignment scores and perform fine-grained evaluations. We address this by using a vision-language model to directly predict both overall and fine-grained alignment scores for image-text pairs. For fine-grained evaluation, TIFA [16], VQ2 [45], and Gecko [42] split prompts into elements and generate questions to assess fine-grained issues in generated images using VQA models. However, these generated questions can sometimes diverge from the original prompt’s content. To address this, we design a new positive-negative question template that provides VQA models with more accurate prompt context, improving alignment with human preference.
@@ 3 EvalMuse-40K Benchmark
In this section, we detail EvalMuse-40K, a reliable and fine-grained benchmark with comprehensive human annotations for T2I evaluation. We present the overview construction of EvalMuse-40K in Fig. 1. In the construction process, we employ various strategies to ensure the diversity and reliability of our benchmark. Next, we introduce the construction process of EvalMuse-40K from three aspects: prompt collection, image generation, and data annotation followed by a statistical analysis of the data.
@@ 3.1 Prompts Collection
Real Prompts from Community. Most prompts used by current benchmarks are generated from templates with LLMs or manually crafted, with only a small portion originating from real-world users. This results in a gap between model evaluation and actual user needs. To address this, we sample 2K real prompts from DiffusionDB [41], which contains 1.8M prompts specified by real users.
To be specific, we first randomly sample 100K prompts from DiffusionDB as the source dataset. To ensure diversity in the final sampled prompts, we classify the prompts in four aspects: subject category, logical relationship, image style, and BERT [18] embedding score. The specific categories of each aspect are shown in the supplementary material. Except for the BERT embedding score, which is computed, the other aspects are categorized using GPT-4 [1], and a prompt can belong to more than one category. After labeling the 100K prompts, we improve the data shaping method proposed in [39] by using Mixed Integer Linear Programming (MILP) to ensure that the sampled data is approximately uniformly distributed across each category. The specific sampling strategy and the distribution of the sampled data are shown in the supplementary material.
Synthetic Prompts for Various Skills. To comprehensively evaluate the generative models’ skills, we employ specific templates and a rich corpus to generate 2K synthetic prompts using GPT-4. These prompts are divided into six categories: (1) Object Count; (2) Object Color and Material; (3) Environment and Time Setting; (4) Object Activity and Perspective Attributes; (5) Text Rendering; and (6) Spatial Composition Attributes. For each category, we use different templates and GPT-4 to generate reasonable and natural prompts. The specific generation methods and details are provided in the supplementary material.
Element Splitting and Question Generation. To achieve fine-grained annotation and evaluation, we perform element splitting and question generation on the 4K collected prompts. In contrast to the word-level annotation used in RichHF [25] and Gecko [42], we split the prompt into fine-grained elements, which increases the consistency of the annotations across different humans. In addition, categorizing each element allows us to examine the model’s capabilities in various aspects at a fine-grained level. We adopt the element categorization strategy from TIFA [16] but refine the question generation process by dividing it into two steps: first, splitting the prompts into individual elements, and then generating specific questions for each element. We generate only simple judgment questions with yes or no answers, making the question generation process more controllable and the evaluation more explainable. The generated questions are also filtered and regenerated to ensure that each element has a corresponding question. The templates used for element splitting and question generation are shown in the supplementary material.
@@ 3.2 Image Generation
To ensure the diversity of the generated images, we select 20 different types and versions of diffusion-based generative models, considering the strong performance of diffusion models [14]. Meanwhile, to ensure the differentiation among the selected models, we classify the selected models into four groups: (1) basic stable diffusion [33] models, such as SD v1.2, SD v1.5, SD v2.1; (2) advanced open-source generative models, such as SDXL [31], SSD1B [11], SD3 [10], HunyuanDiT [24], Kolors [38], PixArt-$\alpha$ [6], PixArt-$\Sigma$ [4], IF [8], Kandinsky3 [2], Playground v2.5 [22]; (3) efficient generative models, such as SDXL-Turbo [36], LCM-SDXL, LCM-SSD1B [29], LCM-PixArt [5], SDXL-Lightning [26]; and (4) proprietary generative models, such as Dreamina [9] and Midjourney v6.1 [15]. For each prompt, we randomly sample a subset of models for image generation, applying default parameters to ensure the quality of the generated images. This process results in 40K image-text pairs being generated using 4K prompts, ensuring a diverse dataset for annotation.
@@ 3.3 Data Annotation
In this section, we describe how we perform the data annotation. We first define the content and templates of the annotations and then detail the entire annotation process.
Annotation Content and Templates. The annotation includes image-text alignment, structural problems, and extra information (see Fig. 1(b)). In terms of image-text alignment, the annotator first scores the alignment using a 5-point Likert scale, like TIFA [16]. Then, for the fine-grained elements in the prompt, the annotator needs to label whether they are aligned with the image or not. Annotators also mark structural issues in the generated images, categorized into three main types: humans, animals, and objects. For human figures, structural issues are further subdivided by specific regions, such as the face, hands, and limbs. To address potential inaccuracies in element splitting by GPT-4, we introduce a new splitting confidence label, enabling annotators to flag instances with incorrect splitting during annotation. Additionally, because some prompts originate from real users and may contain unclear meanings, we add a label to indicate whether a prompt is meaningful.
Annotation Process. To improve the quality of annotation, our annotation process is divided into three stages. (1) Pre-annotation: We formulate straightforward annotation standards to train the annotators, using a small amount of data for pre-annotation. The pre-annotated data are then carefully reviewed. Based on the issues encountered during the annotation and review process, we refine the evaluation standards to ensure greater clarity and consistency. (2) Formal annotation: During formal annotation, each image-text pair is labeled by three annotators, with an additional annotator assigned for quality control. Furthermore, the annotators will identify any NSFW content in the generated images and determine whether they should be discarded. (3) Re-annotation: For instances where alignment scores from the three annotators show significant discrepancies (range $\geq 2$), we conduct re-annotation to reduce subjective bias. Ultimately, each image-text pair is annotated by 3 to 6 annotators, and the average score is used as the final label.